OpenStudy (anonymous):

A rectangle has a length of 9.00 inches and a width of 4.25 inches. What is the area in square inches? Round to the correct significant digits.

OpenStudy (faiqraees):

Area = 9*4.25 Calculate it and then round it to three significant digits

OpenStudy (anonymous):

how do u do that

OpenStudy (janu16):

A=wl=4.25·9=38.25in² @Fallingstar80

OpenStudy (janu16):

so its 38.25in^2

OpenStudy (faiqraees):

@Janu16 They also want to round it to correct significant digits. So it has to be rounded to 3 significant digits.

OpenStudy (janu16):

she can atleast figure tht out by herself

OpenStudy (anonymous):

whats significant digits.

OpenStudy (jay-bird):

signification figure/digits is a way of simplifying numbers, in a way. in this number 6.00, there is 1 significant digit, 6, because .00 has no effect on the number, however in 6.09, there are 3 significant digits because 0 is holding the place of tenths, and 6 and 9 are because they are numbers in places.